OpenStudy (anonymous):

help! The Flat Rock auto assembly plant in Detroit, Michigan, produces three different makes of automobiles. In two years, the plant constructed a total of 390,000 cars. If 90,000 more cars were made in the first year than in the second year, how many cars were made in each year?

OpenStudy (paxpolaris):

let's say they made x in the first year and y in the second year...

OpenStudy (paxpolaris):

\[\large x+y=390,000\] ...correct?

OpenStudy (paxpolaris):

and we also know:\[x=90,000+y\] substitute this in the first equation:\[(90,000+y)+y=390,000\] and then solve for y ... and then get x

OpenStudy (anonymous):

so x would be 180,000, right?

OpenStudy (anonymous):

Right

OpenStudy (paxpolaris):

\[90,000+2y=390,000\]

OpenStudy (paxpolaris):

\[2y=300,000\]

OpenStudy (paxpolaris):

\[y=150,000\]

OpenStudy (paxpolaris):

\[x=90,000+150,000\]

OpenStudy (paxpolaris):

\[x=240,000\]